Find the modulus and the principal argument of each of the given complex numbers. (a) 3 −4i, (b) −2+i, (c) 1 1 +i√ 3 , (d) 7 −i −4−3i, (e) 5(cosπ/3 + isinπ/3), (f) cos2π/3−sin2π/3.
1
Expert's answer
2020-05-28T18:28:16-0400

The formula for modulus and principle argument are as follows

"| z | = \sqrt{\smash[b]{x^2 + y^2}}"



"\theta = arctan(y/x)"


A) 3 - 4i


"| 3 - 4i | = \sqrt{\smash[b]{(3)^2+ (-4)^2}}"

r = "\sqrt{\smash[b]{9+ 16}}"

r = "\sqrt{\smash[b]{25}}"

r = 5


"\theta = arctan(-4/3)"

"\theta" = - arctan(4/3)

"\theta" = - 53.13°



B) -2 + i


"| -2 + i | = \sqrt{\smash[b]{(-2)^2 + (1)^2}}"

"r = \sqrt{\smash[b]{4+ 1}}"

"r = \sqrt{\smash[b]{5}}"


"\arctan(1/-2)"

"= - \arctan(1/2) = -26.56°"

Since the given complex number lies in the second quadrant.

Therefore, "\theta =( -26.56 + 180)°"

"\theta = 153.44°"


C)"z = 1/(1 + i\sqrt{\smash[b]{3}})"

"z = (1 - i\sqrt{\smash[b]{3}})/((1 + i\sqrt{\smash[b]{3}})(1 - i\sqrt{\smash[b]{3}}))"

"z = (1 - i\sqrt{\smash[b]{3}})/(1² - (i\sqrt{\smash[b]{3}})²)"

"z = (1 - i\sqrt{\smash[b]{3}})/(1 + 9)"

"z = (1 - i\sqrt{\smash[b]{3}})/10"

"z = (1/10) - i\sqrt{\smash[b]{3}}/10"


"Modulus, r = \sqrt{\smash[b]{(1/10)² + (\sqrt{\smash[b]{3}}/10)²}}"

"r = \sqrt{\smash[b]{(1/100) + (3/100}}"

"r = \sqrt{\smash[b]{(4/100)}}"

"r = 2\/10"


"\theta = arctan( -(\sqrt{\smash[b]{3}}/10)/(1/10)"

"\theta = arctan(- \sqrt{\smash[b]{3}})"

"\theta = - ( \pi / 3)"

"\theta = - 60°"



D)"z = (7 - i)\/(-4 - 3i)"

"z = ((7 - i)(-4 + 3i))\/((-4 - 3i)(-4 + 3i))"

"z = (-28 + 21i + 4i + 3)\/((-4)\u00b2 - (3i)\u00b2)"

"z = (-25 + 25i)\/(16 + 9)"

"z = (-25 + 25i)\/25"

"z = -1 + i"


"Modulus, r = \sqrt{\smash[b]{(-1)² + 1²}}"

"r = \sqrt{\smash[b]{1 + 1}}"

"r = \sqrt{\smash[b]{2}}"


"\arctan(1/(-1))"

"=\arctan(-1)"

"= -45\degree"

Since the given complex number lies in the second quadrant,

Therefore, "\theta=(-45 + 180)°"

"\theta = 135°"



E) "5(cos(\pi/3) +isin(\pi/3))"

"=5cos(\pi/3) +i5sin(\pi/3)"


"Modulus, r = \sqrt{\smash[b]{(5cos(\pi/3))² + (5sin(\pi/3))²}}"

"r = \sqrt{\smash[b]{5² (cos²(\pi/3) + sin²(\pi/3))}}"

"r = 5\sqrt{\smash[b]{1}}"

"r = 5"


"\theta = arctan( 5sin(\pi/3) / 5cos(\pi/3) )"

"\theta = arctan(tan(\pi/3))"

"\theta =\pi/3 = 60\degree"


F)"z = cos(2\pi/3) - sin(2\pi/3))"

"z = (-1/2) - (\sqrt{\smash[b]{3}}/2) = -(1 + \sqrt{\smash[b]{3}})/2"


"Modulus, r =|-(1+\sqrt{3})/2|=(1+\sqrt{3})/2"

Since the given complex number is negative real number,

"\theta =180°"